OpenStudy (anonymous):

4. A local pizzeria offers 13 toppings for their pizzas and you can choose any 3 of them for one fixed price. How many different types of pizzas can you order with 3 toppings? (1 point)

OpenStudy (anonymous):

It would be (13 C 3). Do you know how to do combinations?

OpenStudy (anonymous):

Aright with combinations you are choosing r ways from the total n ways. In this case. N is the total amount of toppings. R is the amount that you use each time

OpenStudy (anonymous):

So r is 3. And the formula for nCr is n!/(n-r)!)(r!)

OpenStudy (anonymous):

yes do you know what a factorial (!) is?

OpenStudy (anonymous):

no, my techer explains nothing!:/

OpenStudy (anonymous):

alright. so in this case n! would be 12! 12! is an equivalent of 12x11x10x9x8x7x6x5x4x3x2x1

OpenStudy (anonymous):

and (n-r)! would be (12-3)!=9!

OpenStudy (anonymous):

im still confused.. this is my worst subject :/

OpenStudy (anonymous):

so the 12! in the numerator would cancel with the 9! in the denominator. and become 12x11x10

OpenStudy (anonymous):

so then i would multiply those and get the the answer?

OpenStudy (anonymous):

\[\frac{ 12! }{ (12-3)!(3!) }\]

OpenStudy (anonymous):

so the 12! is 12x11x10x9x8x7x6x5x4\[\frac{ 12x11x10x9x8x7x6x5x4x3x2x1 }{ 9x8x7x6x5x4x3x2x1 (3!)}\]

OpenStudy (anonymous):

now do you see that everything from 9 down can cancel out?

OpenStudy (anonymous):

This would equal to:\[\frac{ 12x11x10 }{ 3x2x1 }\]

OpenStudy (anonymous):

o okay

OpenStudy (anonymous):

Are you taking AP statistics?

OpenStudy (anonymous):

no,algebra

OpenStudy (anonymous):

Wow i learned this in statistics. Does it make sense now?

OpenStudy (anonymous):

a little but not really .. lets just say i dont think I could pass the test or come any where near :/

OpenStudy (anonymous):

if you want make me a fan...i need to go.

OpenStudy (anonymous):

this is what i have to chose form 87,178,291,200 14 1 17,297,280

OpenStudy (anonymous):

12*

OpenStudy (anonymous):

I got 220...

OpenStudy (anonymous):

ok thank you for your help I will try and figuer this out :/

OpenStudy (anonymous):

Alright one more thing. I used combinations because order does not matter. This means that ABC are the same as ACB. So a pizza with 3 toppings is the same as a pizza with the same 3 toppings spread around differently

OpenStudy (anonymous):

okay thanks :)

OpenStudy (anonymous):

Sure and i can't believe that in algebra one permutations and combinations are taught. Well i g2g cya!

OpenStudy (anonymous):

hey i answerd 286 and got it correct :)

OpenStudy (anonymous):

My bad it was 13 as n so it would be...\[\frac{ 13x12x11x }{ 3x2x1 }\]

OpenStudy (anonymous):

so \[\frac{ 13! }{ 10!(3!) }\]

OpenStudy (anonymous):

anyway cya i really g2g